A Bibliography of The X Journal: Computing Technology with the X Window System

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

11 April 2012 Version 1.04

Abstract

This bibliography records the contents of the monthly journal The X Journal: Computing Technology with the X Window System, ISSN 1056-7003, CODEN XJOUEA.

Correspondence can be addressed to:

SIGS Publications 71 West 23rd Street, 3rd floor New York, NY 10010 USA

WWW:  http://www.sigs.com/
publications/docs/txjr/

Title word cross-reference

3GX [Mut96b].

5.3 [Aro96b].

Absolute [KF96]. achieve [Cah96]. action [Bru95c]. Ad [Ano96a]. After [Dic96]. alive [Bow96e]. already [HM95]. always [Dic96]. Any [Ran96c]. Application [Ber96b, Lee95, Red96, Bor95, Cah96]. Applications [Skl95b, HM95, SZ95, You95]. Architectural [Log96b]. area [SZ95]. Automated [Le 96]. available [HM95]. Awards [Bru95a].

Back [Bow96f]. Based [Ber96a, Aud95]. Basic [Tro96a]. battle [Dic96]. be [Bor95, Dic96, Kil95]. been [Sch95]. best [Dic96]. Book [Str96]. box [Skl95a]. Brevity [SS96b]. Broadway [Bow96e, Dic96, Lav96]. Browsing [Ran96a, Ran96b]. busy [SZ95].

C [Ber96a, Ber96b, Fur96a, Luk96]. Callbacks [Fur96a]. Can [Le 96, SZ95]. capabilities [Lav96]. capability [KF96]. CDE [Sch95, Und95]. Center [Bru95c]. Choice [Bru95a]. Choose [Ran96c]. Clean [Ran96a, Ran96b]. Client [Min96b, New96, Pro96]. Clients [Bud96, Min96b]. color [Kil96c]. Common [SZ95, Sch95, Und95]. Communicating [Bud96]. complete [SZ95]. Component [Ran96a]. Components [Red96]. computing [Whi95]. concentrate [SZ95]. Concerns [Spe96a]. configurations [Bor95, Und95]. Configuring [Spe96b]. conflicting [Bor95]. Connection [Era96]. CORBA [HM95]. Count [Le 96]. County [Bru95c]. Creating [Log96a, Min96b]. Criteria [Luk96]. Cross [Fur96a, Fur96b]. Cross-Platform [Fur96b]. Currents [Fur96a, Fur96b]. customization [Bor95].

Dane [Bru95c]. Database [Min96a]. debugging [Lee95, You95]. delivers [Sch95]. depth [Bru95b]. Design [Bil95, SS96a, SS96b, Spe96b]. Desktop [Aud96b, SZ95, Und95, Sch95, SZ95]. Developers [Spe96a]. Developing [Bil96b, Red96]. Development [Ber96a, Luk96]. directories [Bor95]. Dispatch [Bru95c]. display [Aud95]. doesn’t [Dic96].

easy [Bor95]. Editor [Ano96b, Bow96a, Bow96b, Bow96c, Bow96d, Red96, Bru95a].

Emacs [Ber96a]. Embedding [Skl95b]. Emergency [Bru95c]. engineering [Lee95]. Enhancing [HM95]. enough [Aud95]. enterprise [Whi95]. entrant [Ber96b]. environment [Sch95, Whi95, SZ95, Und95]. Envizex [Bro96a]. Establishing [Era96]. Evaluation [Luk96]. evolution [Aud95]. Examining [Nan96]. executing [Log96a]. eXit [Whi95]. exotic [Bor95]. expands [KF96]. Extending [Pro96]. extension [KF96].

FAQs [Lew95, Lew96d, Lew96a, Lew96b, Lew96c]. fast [Aud95]. Fat [New96]. favorite [LLM^{+}96]. Files [ALM^{+}96]. flexible [Bor95]. Fly [Fet96b]. follow [Lav96]. follow-on [Lav96]. Font [Ran96c]. Free [Kil96b, Tro96d, Tro96a, Tro96c, Tro96b]. Frontier [Aud96a]. fulfill [Bor95]. Future [Bow96e, Bow96f, Pro96].

general [Bor95]. Generalized [Fur96a]. Generating [Fet96b]. Generation [Dic96]. gets [Bru95c]. GIFs [Fet96b]. GMS [Aro96b]. Gnu [Ber96a]. goal [SZ95]. goals [Bor95]. Graphical [Aro96b, Ran96d, Spe96b]. Graphics [Skl95b, SZ95]. Group [KF96]. GUI [SS96a, SS96b, Spe96a, Spe96b]. Gymnastics [Ran96e].

happenings [Win95]. Heart [Aud96b].

Hello [Bow95].

Highlights [Gil96b, Gil96c, Gil96d, Bru95a]. hooked [Bru95c]. Human [Bil96b, Bil96a].

Implementation [Spe96b]. implementations [HM95]. independence [Sch95]. Index [Ano96a]. Indigo [SZ95]. industrial [HM95]. Industry [Bru95a, Gil96b, Gil96c, Gil96d, Gil96a]. infinite [Und95]. Information [SW96]. input [Lee95].

Insure [Aro96a]. integrated [SZ95]. Integration [Aud96a]. Interface [Spe96b]. Interfaces [Bil96b, Bil96a]. Internetwork [Aud96b]. Internetworking [Aud96a, Aud96b]. interoperability [Sch95]. Interview [Bow96f]. Investment [Lew96d].

Java [Lav96]. Just [Lew95, Lew96d, Lew96a, Lew96b, Lew96c].

key [Dic96].

Land [Ber96a, Ber96b]. Language [Tro96d, Lew95]. layer [KF96]. Layering [KF96]. Lean [Ran96a, Ran96b]. Let [Kil95]. Letter [Red96]. Letters [Ano96b]. Light [Lav96, Kil95]. Linux [Tro96b]. List [Ano96c]. living [Bow96e, Lew95].

Machine [Ran96b]. Magic [SZ95, SZ95]. maintain [Bor95]. marketplace [Dic96]. Maximizing [Kil96c]. Medical [Red96]. message [Skl95a]. Meta [Tro96c]. method [Bor95]. Modeling [Aro96b]. Modes [Ber96b]. Monitors [Min96b]. Motif [Nan96, Skl95a, Win96]. Mousing [SS96a]. MOX [KF96]. Multimedia [Fet96a, Fet96b]. Multiple [KF96]. must [Bor95]. my [Cah96].

nearly [Und95]. News [Gil96a, Gil96e, Gil96f, Gil96g, Bru95b]. nonstandard [Bor95]. NT [Aud96a].

Object [Bil95, Bil96a]. Object-Oriented [Bil96a, Bil95]. Objects [Bil96b]. OnNet [Bro96b]. OOX [Bil96b]. OpenGL [Kil95, Kil96b, Kil96c, Kil96a]. Options [Spe96a]. Oriented [Bil96a, Bil95]. Our [LLM^{+}96]. over/underlay [KF96]. overlay [KF96]. overriding [SZ95]. Overview [Log96b, Und95]. own [Lav96].

Pages [Skl95b]. Paradigm [Pro96]. Parent [Era96]. Parent/Child [Era96]. Part [Ran96b, Ran96e]. patterns [Bil95]. PC [Aud95, Aud96a, Aud96b, Pro96]. PC-based [Aud95]. Child [Era96]. Motif [Skl95a]. people [Lav96]. Perl [Ran96d]. perspective [Whi95]. photoflos [Log96a]. plasticae [SS96a]. Platform [Fur96b, Sch95]. populating [Log96a]. Portability [Spe96a]. powerful [Bor95, Lav96]. Practical [Skl95a]. Predictions [ALM^{+}96]. primer [Bow95]. Product [Aro96a, Aro96b, Bro96a, Bro96b, Bru95b, Gil96a, Gil96e, Gil96f, Gil96g, Mut96b, Mut96a]. Products [Ano96c, LLM^{+}96]. professionals [SZ95]. Programming [Log96a, Log96b, Tro96a]. projects [HM95]. Prototyping [Bil96a]. proven [HM95]. provide [SZ95]. Python [Tro96d, Tro96a, Tro96c].

quality [Cah96].

Radical [Skl95b]. Rapid [Bil96a]. rather [SZ95]. Re [Ber96b]. Re-entrant [Ber96b]. Related [Ano96c]. Relations [Min96a]. reminds [Lav96]. Representing [Min96a]. require [Bor95]. required [Cah96]. resolution [Kil96c]. resource [Bor95]. Reuse [Win96]. Review [Aro96a, Aro96b, Bro96a, Bro96b, Mut96b, Mut96a, Str96]. Revolution [Ran96c, Ran96d, Ran96e, Whi95]. Rework [Win96]. right [Lav96]. Robert [Bow96f]. rodentia [SS96a]. Roundup [LLM^{+}96].

Saving [Lew96d]. Scheifler [Bow96f]. Screen [SS96b]. Scripting [Ran96c, Ran96d, Ran96e].

Secrets [Skl95a]. Selected [Ano96c]. Series [Bro96a]. Server [Spe96b, KF96]. Servers [Aud96b, Pro96, Wal96, Aud95]. set [Lav96]. sets [Bor95, KF96]. Several [HM95, Bor95]. Shedding [Lav96]. should [Cah96]. Silicon [SZ95]. simple [Bor95]. SL-GMS [Aro96b]. Sockets [Bud96]. Software [Fur96b, Kil96b, Lew96d, Tro96d, Tro96a, Tro96c, Tro96b, Cah96, HM95, Lee95]. Solaris [Mut96a]. Solution [Fur96b]. Solutions [Tro96d, Tro96a, Tro96c, Tro96b]. some [HM95]. soul [SS96b]. SPARCbook [Mut96b]. Stations [Wal96, Bro96a]. strategy [Dic96]. style [Bow95]. Styling [Bor95]. supports [KF96, Und95]. survival [Dic96]. Syntactic [Ran96e]. System [Aro96b]. Systems [SW96].

technology [Dic96, Und95]. Tektronix [Bru95b]. test [Cah96]. Tests [Le 96]. their [SZ95]. them [HM95]. themselves [HM95]. there [Kil95]. Thin [New96, Pro96]. Today [Aud96b, Spe96a]. Tool [Ber96a]. Tools [Luk96, You95]. Tree [Min96a]. two [KF96]. types [Und95]. Tyrannosaurus [SS96a].

underlay [KF96]. unified [Sch95]. UNIX [Aud96a, Sch95, SZ95, Whi95]. UNIX/X [Aud96a]. User [Spe96b, Lew96a]. using [SZ95].

waiting [Sch95]. Web [Fur96b, Ran96a, Ran96b, Skl95b]. Web-browsing [Ran96b]. well [Bow96e]. While [Lav96]. Widgets [Tro96c]. will [Dic96]. win [Dic96]. WinDD [Bru95b]. Windows [Win95, Aud96a, Win95]. Wisdom [SW96]. wit [SS96b]. work [SZ95]. World [Bow95]. wrestling [SZ95]. Writer [LLM^{+}96, ALM^{+}96]. writing [SS96b]. WWW [Pro96].

X [Ano96c, Aud95, Aud96a, Aud96b, ALM^{+}96, Ber96a, Bil95, Bil96b, Bil96a, Bow95, Bow96e, Bow96f, Bru95c, Dic96, Fet96a, Fet96b, KF96, Lav96, Lee95, Lew95, Lew96d, Lew96a, Min96b, Pro96, Skl95b, Spe96b, Tro96a, Wal96, You95]. X-Based [Ber96a]. X-Related [Ano96c]. X11 [HM95]. x86 [Mut96a]. Xcetera [You95]. XIE [Log96a, Log96b]. XIElib [Log96a, Log96b]. XT [Skl95a, Bud96]. XT/Motif [Skl95a].

year [Win95]. you’ve [Sch95].

References

Auditore:1996:WP

[ALM^{+}96] Peter Auditore, David B. Lewis, Tom Munnecke, Scott Raney, Mark Kilgard, Erick Schaffer, John Sorflaten, and Daniel Bernstein. Writers’ predictions: From the X files. The X Journal: Computing Technology with the X Window System, 5(5):44, 48, May 1996. CODEN XJOUEA. ISSN 1056-7003.

Anonymous:1996:AI

[Ano96a] Anonymous. Ad index. The X Journal: Computing Technology with the X Window System, 6(2):49, November and December 1996. CODEN XJOUEA. ISSN 1056-7003. URL http:

                                                  //www.sigs.com/publications/
                                                  docs/txjr/9611/txjr9611.toc.
                                                  html.

Anonymous:1996:LE

[Ano96b] Anonymous. Letters to the editor.

The X Journal: Computing Technology with the X Window System, 6(2):6–??, November and December 1996. CODEN XJOUEA. ISSN 1056-7003. URL http:

//www.sigs.com/publications/
docs/txjr/9611/txjr9611.toc.
html.

Anonymous:1996:SLX

[Ano96c] Anonymous. A selected list of 200 X-related products. The X Journal: Computing Technology with the X Window System, 5(6):32– ??, June 1996. CODEN XJOUEA. ISSN 1056-7003.

Arora:1996:PRI

[Aro96a] Ashish Arora. Product review: Insure++. The X Journal: Computing Technology with the X Window System, 5(6):55–??, June 1996. CODEN XJOUEA. ISSN 1056-7003.

Arora:1996:PRS

[Aro96b] Ashish Arora. Product review:

SL-GMS (Graphical Modeling System) 5.3. The X Journal: Computing Technology with the X Window System, 5(5):90–??, May 1996. CODEN XJOUEA. ISSN 1056-

7003. URL http://www.sigs.
com/publications/docs/txjr/
9605/txjr9605.d.prod.html.

Auditore:1995:PXE

[Aud95] Peter Auditore. PC-based X: The evolution of PC X display servers: How fast is fast enough? The X Journal: Computing Technology with the X Window System, 4 (3):67–??, January/February 1995. CODEN XJOUEA. ISSN 1056-7003.

Auditore:1996:PIN

[Aud96a] Peter Auditore. PC internetworking: A new frontier — Windows NT and UNIX/X Windows integration. The X Journal: Computing Technology with the X Window System, 5(5):70–??, May 1996. CODEN XJOUEA. ISSN 1056-7003.

Auditore:1996:PIP

[Aud96b] Peter Auditore. PC internetworking: PC X servers are the heart of today’s internetwork desktop. The X Journal: Computing Technology with the X Window System, 5(6):18–??, June 1996. CODEN XJOUEA. ISSN 1056-7003.

Bernstein:1996:LCG

[Ber96a] Daniel J. Bernstein. From the land of C++: Gnu Emacs: An X-based development tool. The X Journal: Computing Technology with the X Window System, 5(6):77–??, June 1996. CODEN XJOUEA. ISSN 1056-7003.

Bernstein:1996:LCR

[Ber96b] Daniel J. Bernstein. From the land of C++: Re-entrant application modes. The X Journal: Computing Technology with the X Window System, 5(5):80–??, May 1996. CODEN XJOUEA. ISSN 1056-7003.

Bilow:1995:OXD

[Bil95] Steven Bilow. Object-oriented X:

Design patterns. The X Journal: Computing Technology with the X Window System, 4(3):77–??, January/February 1995. CODEN XJOUEA. ISSN 1056-7003.

Bilow:1996:OXR

[Bil96a] Steven C. Bilow. Object-oriented X: Rapid prototyping of human interfaces. The X Journal: Computing Technology with the X Window System, 5(7):??, July/August 1996. CODEN XJOUEA. ISSN 1056-7003.

Bilow:1996:ODH

[Bil96b] Steven C. Bilow. OOX: Developing human interfaces with objects and X. The X Journal: Computing Technology with the X Window System, 5(6):73–??, June 1996. CODEN XJOUEA. ISSN 1056-7003.

Bornetun:1995:SYA

[Bor95] Ulrika Bornetun. Styling your application with resource sets: A general customization method must fulfill several conflicting goals: it has to be powerful, yet simple; it must be flexible, yet easy to maintain; and it must not require exotic configurations or nonstandard directories. The X Journal: Computing Technology with the X Window System, 4(3):50–??, January/February 1995. CODEN XJOUEA. ISSN 1056-7003.

Bowman:1995:XPH

[Bow95] Charles F. Bowman. The X primer: Hello world — X style. The X Journal: Computing Technology with the X Window System, 4(3):70–??, January/February 1995. CODEN XJOUEA. ISSN 1056-7003.

Bowman:1996:Ea

[Bow96a] Charles F. Bowman. From the editor. The X Journal: Computing Technology with the X Window System, 5(5):4, May 1996. CO-DEN XJOUEA. ISSN 1056-7003.

Bowman:1996:Eb

[Bow96b] Charles F. Bowman. From the editor. The X Journal: Computing Technology with the X Window System, 5(6):4, June 1996. CO-DEN XJOUEA. ISSN 1056-7003.

Bowman:1996:Ec

[Bow96c] Charles F. Bowman. From the editor. The X Journal: Computing Technology with the X Window System, 5(7):??, July/August 1996. CODEN XJOUEA. ISSN 1056-7003.

Bowman:1996:Ed

[Bow96d] Charles F. Bowman. From the editor. The X Journal: Computing Technology with the X Window System, 6(2):4–??, November and December 1996. CO-DEN XJOUEA. ISSN 1056-

                                        7003. URL http://www.sigs.
                                        com/publications/docs/txjr/
                                        9611/txjr9611.toc.html.

Bowman:1996:FXX

[Bow96e] Charles F. Bowman. The future of X: X is alive and well (and living on Broadway!). The X Journal: Computing Technology with the X Window System, 5(5):30–??, May 1996. CODEN XJOUEA. ISSN 1056-7003.

Bowman:1996:XBF

[Bow96f] Charles F. Bowman. X: Back to the future: An interview with Robert Scheifler. The X Journal: Computing Technology with the X Window System, 5(5):38–43, May 1996. CODEN XJOUEA. ISSN 1056-7003. Discusses relationship of Broadway to Java.

Brooks:1996:PRE

[Bro96a] Darryl Brooks. Product review:

Envizex I Series Stations. The X Journal: Computing Technology with the X Window System, 5(6):60–??, June 1996. CODEN XJOUEA. ISSN 1056-7003.

Brooks:1996:PRO

[Bro96b] Darryl Brooks. Product review:

OnNet. The X Journal: Computing Technology with the X Window System, 5(6):53–??, June 1996. CODEN XJOUEA. ISSN 1056-7003.

Bruno:1995:IHE

[Bru95a] Lee Bruno. Industry highlights:

Editor’s Choice Awards. The X Journal: Computing Technology with the X Window System, 4 (3):6–??, January/February 1995. CODEN XJOUEA. ISSN 1056-7003.

Bruno:1995:PND

[Bru95b] Lee Bruno. Product news: In depth: Tektronix WinDD. The X Journal: Computing Technology with the X Window System, 4 (3):82–??, January/February 1995. CODEN XJOUEA. ISSN 1056-7003.

Bruno:1995:XAD

[Bru95c] Lee Bruno. X in action: Dane County Emergency Dispatch Center gets hooked on X. The X Journal: Computing Technology with the X Window System, 4(3):89–??, January/February 1995. CODEN XJOUEA. ISSN 1056-7003.

Budne:1996:XCC

[Bud96] Daniel Budne. Xt clients communicating with sockets. The X Journal: Computing Technology with the X Window System, 6 (2):22–??, November and December 1996. CODEN XJOUEA. ISSN 1056-7003. URL http:

//www.sigs.com/publications/
docs/txjr/9611/txjr9611.toc.
html.

Cahill:1996:HST

[Cah96] Conor P. Cahill. How should I test my application to achieve the required software quality? The X Journal: Computing Technology with the X Window System, ??(??):??, ???? 1996. CO-DEN XJOUEA. ISSN 1056-

7003. URL http://www.sigs.
com/publications/docs/txjr/
9603/txjr9603.f.cahill.html.

Dichter:1996:AXG

[Dic96] Carl Dichter. After the X generation: The best technology doesn’t always win the battle of the marketplace — strategy will be key to Broadway — and X’s — survival.

The X Journal: Computing Technology with the X Window System, 5(5):46–??, May 1996. CODEN XJOUEA. ISSN 1056-7003.

Erabelli:1996:EPC

[Era96] Prasad Erabelli. Establishing a parent/child connection. The X Journal: Computing Technology with the X Window System, 5 (6):66–??, June 1996. CODEN XJOUEA. ISSN 1056-7003.

Fetik:1996:MX

[Fet96a] Richard Fetik. Multimedia X. The X Journal: Computing Technology with the X Window System, 5(7): ??, July/August 1996. CODEN XJOUEA. ISSN 1056-7003.

Fetik:1996:MXG

[Fet96b] Richard Fetik. Multimedia X:

Generating GIFs on the fly. The X Journal: Computing Technology with the X Window System, 5(5):86–??, May 1996. CODEN XJOUEA. ISSN 1056-7003.

Furr:1996:CCG

[Fur96a] Steven Furr. Cross currents: Generalized callbacks for C++. The X Journal: Computing Technology with the X Window System, 5(7): ??, July/August 1996. CODEN XJOUEA. ISSN 1056-7003.

Furr:1996:CCW

[Fur96b] Steven Furr. Cross currents: The Web as a cross-platform software solution. The X Journal: Computing Technology with the X Window System, 5(6):12–??, June 1996. CODEN XJOUEA. ISSN 1056-

7003. URL http://www.sigs.

                                        com/publications/docs/txjr/
                                        9605/txjr9605.f.furr.html.

Gill:1996:IPN

[Gil96a] Philip J. Gill. Industry and product news. The X Journal: Computing Technology with the X Window System, 5(5):6–??, May 1996. CODEN XJOUEA. ISSN 1056-7003.

Gill:1996:IHa

[Gil96b] Philip J. Gill. Industry highlights. The X Journal: Computing Technology with the X Window System, 5(6):8–??, June 1996. CO-DEN XJOUEA. ISSN 1056-7003.

Gill:1996:IHb

[Gil96c] Philip J. Gill. Industry highlights. The X Journal: Computing Technology with the X Window System, 5(7):??, July/August 1996. CO-DEN XJOUEA. ISSN 1056-7003.

Gill:1996:IHc

[Gil96d] Philip J. Gill. Industry highlights. The X Journal: Computing Technology with the X Window System, 6(2):8–??, November and December 1996. CODEN XJOUEA. ISSN 1056-7003. URL http:

                                        //www.sigs.com/publications/
                                        docs/txjr/9611/txjr9611.toc.
                                        html.

Gill:1996:PNa

[Gil96e] Philip J. Gill. Product news. The X Journal: Computing Technology with the X Window System, 5(6):88–??, June 1996. CODEN XJOUEA. ISSN 1056-7003.

Gill:1996:PNb

[Gil96f] Philip J. Gill. Product news. The X Journal: Computing Technology with the X Window System, 5(7): ??, July/August 1996. CODEN XJOUEA. ISSN 1056-7003.

Gill:1996:PNc

[Gil96g] Philip J. Gill. Product news. The X Journal: Computing Technology with the X Window System, 6 (2):63–??, November and December 1996. CODEN XJOUEA. ISSN 1056-7003. URL http:

//www.sigs.com/publications/
docs/txjr/9611/txjr9611.toc.
html.

Hubert:1995:EXA

[HM95] Richard Hubert and Bernhard Merkle. Enhancing X11 applications with CORBA: Several CORBA implementations are now available and some of them have already proven themselves in industrial software projects. The X Journal: Computing Technology with the X Window System, 4 (3):50–??, January/February 1995. CODEN XJOUEA. ISSN 1056-7003.

Kulesza:1996:MLX

[KF96] John Kulesza and Peter Feil. Multiple layering for X: The multiple over/underlay extension (MOX) to the X server expands X’s overlay capability and supports two layer sets: Absolute and group. The X Journal: Computing Technology with the X Window System, 5(5):52–??, May 1996. CODEN XJOUEA. ISSN 1056-7003.

Kilgard:1995:OLT

[Kil95] Mark Kilgard. OpenGL: Let there be light! The X Journal: Computing Technology with the X Window System, 4(3):12–??, January/February 1995. CODEN XJOUEA. ISSN 1056-7003.

Kilgard:1996:OO

[Kil96a] Mark Kilgard. OpenGL: OpenGL 1.1. The X Journal: Computing Technology with the X Window System, 5(7):??, July/August 1996. CODEN XJOUEA. ISSN 1056-7003.

Kilgard:1996:OFO

[Kil96b] Mark J. Kilgard. OpenGL: Free OpenGL software. The X Journal: Computing Technology with the X Window System, 5(5):74–??, May 1996. CODEN XJOUEA. ISSN 1056-7003.

Kilgard:1996:OMC

[Kil96c] Mark J. Kilgard. OpenGL: Maximizing color resolution. The X Journal: Computing Technology with the X Window System, 6 (2):44–??, November and December 1996. CODEN XJOUEA. ISSN 1056-7003. URL http:

                                        //www.sigs.com/publications/
                                        docs/txjr/9611/txjr9611.toc.
                                        html.

Lavallee:1996:SLB

[Lav96] Paul C. Lavallee. Shedding light on Broadway: While it reminds people of Java, the X follow-on has a set of powerful capabilities in its own right. The X Journal: Computing Technology with the X Window System, 5(5):32–34, 36, May 1996. CODEN XJOUEA. ISSN 1056-7003.

LeSuer:1996:ATY

[Le 96] Brian Le Suer. Automated tests you can count on. The X Journal: Computing Technology with the X Window System, 5(6):62–??, June 1996. CODEN XJOUEA. ISSN 1056-7003. URL http://www.

sigs.com/publications/docs/
txjr/9607/txjr9608.c.lesuer.
html.

Lee:1995:XAS

[Lee95] Kenton Lee. X application software engineering: X input debugging. The X Journal: Computing Technology with the X Window System, 4(3):22–??, January/February 1995. CODEN XJOUEA. ISSN 1056-7003.

Lewis:1995:JFX

[Lew95] David B. Lewis. Just the FAQs: X:

Not a living language. The X Journal: Computing Technology with the X Window System, 4(3):80–??, January/February 1995. CODEN XJOUEA. ISSN 1056-7003.

Lewis:1996:JFXb

[Lew96a] David Lewis. Just the FAQs:

X and the user. The X Journal: Computing Technology with the X Window System, 6(2):12– ??, November and December 1996. CODEN XJOUEA. ISSN 1056-

7003. URL http://www.sigs.
com/publications/docs/txjr/
9611/txjr9611.toc.html.

Lewis:1996:JFa

[Lew96b] David B. Lewis. Just the FAQs. The X Journal: Computing Technology with the X Window System, 5(5):62–??, May 1996. CODEN XJOUEA. ISSN 1056-7003.

Lewis:1996:JFb

[Lew96c] David B. Lewis. Just the FAQs. The X Journal: Computing Technology with the X Window System, 5(7):??, July/August 1996. CO-DEN XJOUEA. ISSN 1056-7003.

Lewis:1996:JFXa

[Lew96d] David B. Lewis. Just the FAQs: X: Saving software investment? The X Journal: Computing Technology with the X Window System, 5(6):96–??, June 1996. CODEN XJOUEA. ISSN 1056-7003.

Anonymous:1996:WRO

[LLM^{+}96] David B. Lewis, Syd Logna, Tom Munnecke, Scott Raney, Erick Schaffer, and John Sorflaten. Writers’ roundup — our favorite products. The X Journal: Computing Technology with the X Window System, 5(6):50–??, June 1996. CODEN XJOUEA. ISSN 1056-7003.

Logan:1996:XPC

[Log96a] Syd Logan. XIElib programming: Creating, populating, and executing XIE photoflos. The X Journal: Computing Technology with the X Window System, 6 (2):50–??, November and December 1996. CODEN XJOUEA. ISSN 1056-7003. URL http:

                                        //www.sigs.com/publications/

docs/txjr/9611/txjr9611.toc.
html.

Logan:1996:XPX

[Log96b] Syd Logan. XIElib programming:

XIE architectural overview. The X Journal: Computing Technology with the X Window System, 5(7): ??, July/August 1996. CODEN XJOUEA. ISSN 1056-7003.

Lukner:1996:ECC

[Luk96] Ralf Lukner. Evaluation criteria for C++ development tools. The X Journal: Computing Technology with the X Window System, 5(6):24–??, June 1996. CODEN XJOUEA. ISSN 1056-7003.

Mintz:1996:RDR

[Min96a] Alex Mintz. Representing database relations as a tree. The X Journal: Computing Technology with the X Window System, 6 (1):40–??, September and October 1996. CODEN XJOUEA. ISSN 1056-7003. URL http://www.

sigs.com/publications/docs/
txjr/9609/txjr9609.f.mintz.
html.

Mintz:1996:CXC

[Min96b] Alexander Mintz. Creating an X client that monitors X clients. The X Journal: Computing Technology with the X Window System, 5(7): ??, July/August 1996. CODEN XJOUEA. ISSN 1056-7003.

Mutreja:1996:PRSb

[Mut96a] Mohit Mutreja. Product review:

Solaris 2.5 x86. The X Journal: Computing Technology with the X Window System, 5(6):58–??, June 1996. CODEN XJOUEA. ISSN 1056-7003.

Mutreja:1996:PRSa

[Mut96b] Mohit Mutreja. Product review: SPARCbook 3GX. The X Journal: Computing Technology with the X Window System, 5(6):51–??, June 1996. CODEN XJOUEA. ISSN 1056-7003.

Nandhra:1996:EM

[Nan96] Ian Nandhra. Examining Motif 2.0. The X Journal: Computing Technology with the X Window System, 5(5):12–??, May 1996. CODEN XJOUEA. ISSN 1056-7003.

Newman:1996:FCT

[New96] Matt Newman. Fat client, thin client. The X Journal: Computing Technology with the X Window System, 5(7):??, July/August 1996. CODEN XJOUEA. ISSN 1056-7003. URL http://www.

                                        sigs.com/publications/docs/
                                        txjr/9607/txjr9607.c.newman.
                                        html.

Prout:1996:FPX

[Pro96] Clive Prout. The future of PC X servers: Extending the WWW thin client paradigm. The X Journal: Computing Technology with the X Window System, 5(7): ??, July/August 1996. CODEN XJOUEA. ISSN 1056-7003.

Rand:1996:LCWa

[Ran96a] Doug Rand. A lean, clean Web browsing component. The X Journal: Computing Technology with the X Window System, 5(3):46– 50, January/February 1996. CO-DEN XJOUEA. ISSN 1056-

7003. URL http://www.sigs.
com/publications/docs/txjr/
9601/txjr9601.f.rand.html.

Rand:1996:LCWb

[Ran96b] Douglas Rand. A lean, clean, Webbrowsing machine: Part 2. The X Journal: Computing Technology with the X Window System, 6(2): ??, November and December 1996. CODEN XJOUEA. ISSN 1056-

7003. URL http://www.sigs.
com/publications/docs/txjr/
9610/txjr9610.f.rand.html.

Raney:1996:SRC

[Ran96c] Scott Raney. The scripting revolution: Choose a font, any font. The X Journal: Computing Technology with the X Window System, 5(7):??, July/August 1996. CO-DEN XJOUEA. ISSN 1056-7003.

Raney:1996:SRG

[Ran96d] Scott Raney. The scripting revolution: Graphical perl. The X Journal: Computing Technology with the X Window System, 6(2):18– 20, November and December 1996. CODEN XJOUEA. ISSN 1056-

7003. URL http://www.sigs.
com/publications/docs/txjr/
9611/txjr9611.raney.html.

Raney:1996:SRS

[Ran96e] Scott Raney. The scripting revolution: Syntactic gymnastics, part 3. The X Journal: Computing Technology with the X Window System, 5(5):64–??, May 1996. CODEN XJOUEA. ISSN 1056-7003.

Reddy:1996:LED

[Red96] Ranthy Reddy. Letter to the editor: Developing a medical application with components. The X Journal: Computing Technology with the X Window System, 5(6):6–??, June 1996. CODEN XJOUEA. ISSN 1056-7003.

Schulman:1995:UDY

[Sch95] Rand R. Schulman. The UNIX desktop you’ve been waiting for: CDE: A common desktop environment for UNIX delivers interoperability, platform independence, and a unified environment. The X Journal: Computing Technology with the X Window System, 4 (3):38–??, January/February 1995. CODEN XJOUEA. ISSN 1056-7003.

Sklar:1995:PXM

[Skl95a] David Sklar. Practical XT/ Motif: Secrets of the Motif message box. The X Journal: Computing Technology with the X Window System, 4(3):75–??, January/February 1995. CODEN XJOUEA. ISSN 1056-7003.

Sklar:1995:RXE

[Skl95b] David F. Sklar. Radical X: Embedding graphics applications inside Web pages. The X Journal: Computing Technology with the X Window System, 5(5):68–69, May 1995. CODEN XJOUEA. ISSN 1056-7003. Discusses Java use on the Web.

Spears:1996:GPC

[Spe96a] Rod Spears. GUI portability: Concerns and options for today’s developers. The X Journal: Computing Technology with the X Window System, 6(2):36–??, November and December 1996. CO-DEN XJOUEA. ISSN 1056-

7003. URL http://www.sigs.
com/publications/docs/txjr/
9611/txjr9611.toc.html.

Spencer:1996:GDD

[Spe96b] Todd Spencer. GUI design — the design and implementation of a graphical user interface for configuring the X server. The X Journal: Computing Technology with the X Window System, 6 (2):26–??, November and December 1996. CODEN XJOUEA. ISSN 1056-7003. URL http:

//www.sigs.com/publications/
docs/txjr/9611/txjr9611.toc.
html.

Schaffer:1996:GDM

[SS96a] Eric Schaffer and John Sorflaten.

GUI design: Mousing around: Tyrannosaurus rodentia plasticae. The X Journal: Computing Technology with the X Window System, 5(5):20–??, May 1996. CODEN XJOUEA. ISSN 1056-7003.

Schaffer:1996:GDS

[SS96b] Eric Schaffer and John Sorflaten.

GUI design: Screen writing: “Brevity is the soul of wit”. The X Journal: Computing Technology with the X Window System, 6 (2):40–??, November and December 1996. CODEN XJOUEA. ISSN 1056-7003. URL http:

//www.sigs.com/publications/
docs/txjr/9611/txjr9611.toc.
html.

Strauss:1996:BR

[Str96] Fred J. Strauss. Book review. The X Journal: Computing Technology with the X Window System, 6 (2):58–??, November and December 1996. CODEN XJOUEA. ISSN 1056-7003. URL http:

                                        //www.sigs.com/publications/
                                        docs/txjr/9611/txjr9611.toc.
                                        html.

Wendell:1996:ISW

[SW96] John Sorflaten and Robert Wendell. Information systems into wisdom systems. The X Journal: Computing Technology with the X Window System, 5(7):??, July/ August 1996. CODEN XJOUEA. ISSN 1056-7003.

Smith:1995:IMC

[SZ95] Amy Smith and Betsy Zeller. Indigo Magic: An (un)Common Desktop Environment: The overriding goal of the Indigo Magic Desktop from Silicon Graphics is to provide an integrated, complete work area so that busy professionals can concentrate on using their applications rather than wrestling with UNIX. The X Journal: Computing Technology with the X Window System, 4 (3):56–??, January/February 1995. CODEN XJOUEA. ISSN 1056-

                                        7003. URL http://www.sigs.
                                        com/publications/docs/txjr/
                                        95/txjr9501s.html.

Troan:1996:FSSb

[Tro96a] Erik Troan. Free software solutions: Basic X programming in Python. The X Journal: Computing Technology with the X Window System, 5(6):84–??, June 1996. CODEN XJOUEA. ISSN 1056-7003.

Troan:1996:FSSd

[Tro96b] Erik Troan. Free software solutions: Linux 2.0. The X Journal: Computing Technology with the X Window System, 6(2):72, 70, November and December 1996. CODEN XJOUEA. ISSN 1056-

7003. URL http://www.sigs.
com/publications/docs/txjr/
9611/txjr9611.toc.html.

Troan:1996:FSSc

[Tro96c] Erik Troan. Free software solutions: Meta widgets in Python. The X Journal: Computing Technology with the X Window System, 5(7):??, July/August 1996. CO-DEN XJOUEA. ISSN 1056-7003.

Troan:1996:FSSa

[Tro96d] Erik Troan. Free software solutions: The Python language. The X Journal: Computing Technology with the X Window System, 5(5):96–??, May 1996. CODEN XJOUEA. ISSN 1056-7003.

Underriner:1995:CTO

[Und95] Kevin R. Underriner. CDE: A technology overview: The Common Desktop Environment supports nearly infinite types of configurations. The X Journal: Computing Technology with the X Window System, 4(3):30–??, January/February 1995. CODEN XJOUEA. ISSN 1056-7003.

Walls:1996:XSX

[Wal96] Jeff Walls. 24 X servers and X stations. The X Journal: Computing Technology with the X Window System, 5(7):??, July/August 1996. CODEN XJOUEA. ISSN 1056-7003.

White:1995:EUR

[Whi95] Gregory White. eXit: The Unix revolution in the enterprise computing environment: A perspective. The X Journal: Computing Technology with the X Window System, 4(3):96–??, January/February 1995. CODEN XJOUEA. ISSN 1056-7003.

Wingo:1995:WHY

[Win95] Scot Wingo. Windows happenings: 1995 — the year of Windows 95? The X Journal: Computing Technology with the X Window System, 4(3):64–??, January/February 1995. CODEN XJOUEA. ISSN 1056-7003.

Winston:1996:RRM

[Win96] Peter Winston. Reuse and rework with Motif. The X Journal: Computing Technology with the X Window System, 5(5):60–??, May 1996. CODEN XJOUEA. ISSN 1056-

                                        7003. URL http://www.sigs.
                                        com/publications/docs/txjr/
                                        9603/txjr9603.f.winston.html.

Young:1995:XTD

[You95] Doug Young. Xcetera: Tools for debugging X applications. The X Journal: Computing Technology with the X Window System, 4 (3):26–??, January/February 1995.

REFERENCES 15 CODEN XJOUEA. ISSN 1056-